# ReMoD: Rethinking Modality Contribution in Multimodal Stance Detection via Dual Reasoning

### v1

@@ Abstract
Multimodal Stance Detection (MSD) is a crucial task for understanding public opinion on social media. Existing work simply fuses information from various modalities to learn stance representations, overlooking the varying contributions of stance expression from different modalities. Therefore, stance misunderstanding noises may be drawn into the stance learning process due to the risk of learning errors by rough modality combination. To address this, we get inspiration from the dual-process theory of human cognition and propose ReMoD, a framework that Rethinks Modality contribution of stance expression through a Dual-reasoning paradigm. ReMoD integrates experience-driven intuitive reasoning to capture initial stance cues with deliberate reflective reasoning to adjust for modality biases, refine stance judgments, and thereby dynamically weight modality contributions based on their actual expressive power for the target stance. Specifically, the intuitive stage queries the Modality Experience Pool (MEP) and Semantic Experience Pool (SEP) to form an initial stance hypothesis, prioritizing historically impactful modalities. This hypothesis is then refined in the reflective stage via two reasoning chains: Modality-CoT updates MEP with adaptive fusion strategies to amplify relevant modalities, while Semantic-CoT refines SEP with deeper contextual insights of stance semantics. These dual experience structures are continuously refined during training and recalled at inference to guide robust and context-aware stance decisions. Extensive experiments on the public MMSD benchmark demonstrate that our ReMoD significantly outperforms most baseline models and exhibits strong generalization capabilities.
Introduction
@@
The proliferation of social platforms supporting multimodal content, such as text or images (8-dar2025social; 9-bansal2024hybrid), has made Multimodal Stance Detection (MSD) a critical task for understanding public opinion (1-liang2024multi; 2-niu2024multimodal). MSD aims to identify the expressed attitude toward a specific target by integrating both textual and visual information, enabling a more comprehensive analysis than purely textual approaches (5-augenstein2016stance; 6-liang2022zero). This task applies widely to political discourse analysis, misinformation detection, and public health communication, making it a crucial tool for decoding complex online interactions (2-niu2024multimodal; 3-pangtey2025large).
Figure 1: Examples of different samples with various modality contributions.
Existing MSD methods focus on enhancing the capability of extracting multimodal representations (1-liang2024multi; 31-zhang2025logic). However, these methods simply fuse information from different modalities to learn stance representations, overlooking the fact that the contributions of different modalities in some examples to the final stance decision can vary. Figure 1 demonstrates how modality contributions vary, illustrating two issues not considered in existing MSD tasks. First, the informative value of individual modalities is highly context-dependent, as seen in Figure 1(a) and (b), where a modality may be ambiguous or irrelevant in isolation, providing little evidential support for stance detection. Models that fail to determine modal contributions are thus susceptible to being skewed by such low-quality inputs. Second, the semantic relationship between modalities is not fixed but dynamic, ranging from complementary to contradictory as shown in Figure 1(c) and (d). Simple feature alignment is ill-equipped to navigate these nuanced interactions. Therefore, when encountering modalities of varying contributions, these methods tend to inappropriately fuse low-quality modal information with high-confidence judgments, thereby contaminating the entire detection process.
To bridge this gap, we introduce ReMoD, a novel framework for multimodal stance detection that Rethinks Modality contribution of stance expression by integrating a Dual-reasoning paradigm inspired by the dual-process theory of human cognition. ReMoD integrates two core reasoning stages to dynamically weight modality contributions based on their actual expressive power for the target stance. First, an experience-driven intuitive reasoning stage instantly generates an initial stance hypothesis by querying the Modality Experience Pool (MEP) and Semantic Experience Pool (SEP), prioritizing historically impactful modality. This initial hypothesis is then passed to a deliberate reflective reasoning stage for critical scrutiny via two complementary reasoning chains. The Modality Chain of Thought (Modality-CoT) reflects on the initial prediction to formulate adaptive fusion strategies, dynamically learning how to weigh different modalities. These strategies are then used to update the MEP, creating a self-improving loop for modality handling. Concurrently, the Semantic Chain of Thought (Semantic-CoT) analyzes the content’s contextual nuances to distill deeper contextual insights, which are updated to SEP. During training, ReMoD continuously refines these dual experience pools, and at inference, it recalls this accumulated experience to make robust, context-aware decisions. Our main contributions are as follows:
  • We propose ReMoD, a novel framework that pioneers a dual-reasoning paradigm for multimodal stance detection. By synergizing experience-driven intuitive reasoning with deliberate reflective reasoning, it realizes a human-like adaptive reasoning process tailored to complex multimodal scenarios, addressing the oversight of dynamic modality contributions in existing methods.
  • We design the Modality Experience Pool and Semantic Experience Pool integrated with a modality-synergistic experience retrieval, which dynamically accumulates modality contribution patterns and semantic insights.
@@
  • Extensive experiments conducted on the public MMSD benchmark demonstrate the superiority of our approach. ReMoD significantly outperforms most baseline models and exhibits strong generalization capabilities.
Figure 2: Overflow of our ReMoD framework for multimodal stance detection.
Related Works
LLM-based Stance Detection
@@
The design of modern stance detection frameworks has been significantly influenced by LLMs. Early approaches focused on harnessing their zero-shot and few-shot capabilities, using direct prompting and Chain-of-Thought (CoT) to elicit stance predictions without labeled data or fine-tuning (13-zhang2023investigating; 14-zhang2022would). Other works have pursued fine-tuning, for instance, by fusing retrieved background knowledge with input pairs to improve learning through backpropagation (11-li2023stance). Subsequent research has sought to further unleash LLM reasoning. Innovations include integrating multi-perspective insights via CoT and prompt-tuning (15-ding2024cross), formalizing reasoning with first-order logic rules for enhanced interpretability (16-zhang2024knowledge), and architecting multi-agent systems where LLMs collaborate for detection (12-lan2024stance; 19-wang2024deem). However, these methods are primarily developed for pure textual stance detection and have not yet explored the integration of multimodal information such as images.
Multimodal Stance Detection
@@
The proliferation of social media and video platforms has driven the rise of multimodal content (text, image) in forms like memes and reels (8-dar2025social; 9-bansal2024hybrid), which convey implicit opinions through diverse signals, posing challenges for stance detection. Modeling semantically heterogeneous cross-modal interactions remains difficult for traditional models, whereas transformer architectures and LLMs excel here via self-attention and transfer learning. Recent approaches increasingly adopt multimodal fusion strategies to improve stance classification. For example, 2-niu2024multimodal proposed MmMtCSD and MLLM-SD for text-image pairs in multi-turn dialogues; 23-kuo2024advancing fused textual, visual, and social signals in political discourse using S-BERT, BEiT, and GraphSAGE. SERN (24-xie2021sern) enhanced fake news detection via sentence-guided visual attention; CT-TN (25-khiabani2023few) and 26-barel2025acquired addressed few-shot/noisy settings with graph-based cues; 27-yao2023end extended MSD to fact-checking via web evidence integration. However, these approaches typically employ rigid fusion strategies, overlooking the dynamic interplay and context-dependent reliability of modalities, which makes them vulnerable to noisy or contradictory information.
Method
@@
This section introduces our novel ReMoD framework for MSD in detail. Drawing inspiration from the dual-process theory of human cognition (47-kahneman2011fast), a psychological model that posits two distinct modes of thought. One is an intuitive, fast ‘System 1’ that relies on heuristics and prior experiences to make rapid judgments. The other is a deliberate, slow ‘System 2’ that engages in analytical reasoning to refine and validate these initial assessments. Our ReMoD framework operationalizes this cognitive paradigm for multimodal stance detection.
@@
Given a text-image pair $(u,v)$, the aim of ReMoD is to identify the stance label $y$, towards a specific target $t$. The architecture of our ReMoD is depicted in Figure 2, which consists of three primary components: 1) Multimodal Knowledge Perception, which extracts comprehensive multimodal information and captures relevant external knowledge. 2) Experience-Driven Intuitive Reasoning (mirroring System 1), which queries the Modality Experience Pool to instantly form an initial stance hypothesis; and 3) Deliberate Reflective Reasoning (emulating System 2), which employs complementary reasoning chains to scrutinize the initial hypothesis and continuously refine the dual experience pools.
Multimodal Knowledge Perception
@@
Textual Knowledge Attainment.
For the text input $u$, we perform a three-step knowledge extraction process. First, we employ a Multimodal Large Language Model (MLLM) with strategically designed prompts to identify a set of key semantic entities, denoted as $\mathcal{E}_{u}$. To ground these entities in external world knowledge, we retrieve corresponding background information for each entity $e\in\mathcal{E}_{u}$ from the Wikipedia knowledge base. As the raw retrieved knowledge is often verbose and contains noisy information, we leverage the MLLM again to distill it into a concise, semantically-rich summary. This final summary, denoted as the text knowledge $\mathcal{K}_{u}$, serves as the knowledge for the text modality.
Visual Knowledge and Feature Grounding.
For the image input $v$, we capture both high-level semantic concepts and low-level visual information. First, to extract semantic concepts, an MLLM generates an image caption, identifies a set of visual entities $\mathcal{E}_{v}$ from the caption. These entities are then used to query Wikipedia, and the resulting information is summarized into a condensed visual knowledge $\mathcal{K}_{v}$. However, knowledge alone is insufficient to capture the visual manifestation of these concepts. To address this, we ground the semantic entities in the image’s pixel space. We leverage the identified entity set $\mathcal{E}_{v}$ to guide the open-vocabulary segmentation model SAM (29-ravi2024sam) in precisely segmenting the image regions corresponding to each entity, yielding a set of entity-aligned visual information $\mathcal{V}$ where each element is a segmented image directly showcasing the visual appearance of a specific semantic entity.
@@
Experience-Driven Intuitive Reasoning
This stage operationalizes the ‘fast thinking’ of our dual-reasoning paradigm. It simulates rapid, intuitive judgment by leveraging experience pools accumulated during training to quickly formulate an initial stance hypothesis. The process consists of two main steps: modality-synergistic experience retrieval and experience-driven stance prediction.
Modality-Synergistic Experience Retrieval.
To retrieve relevant past experiences, we introduce a modality-synergistic retrieval mechanism. Unlike conventional text-only systems, our approach leverages both text and visual information to query the modality experience pool and semantic experience pool. Each pool stores a collection of experiences $\{E_{j}\}$, where each experience is a key-value pair. The key is a pre-computed bimodal embedding $(k_{u}^{j},k_{\mathcal{V}}^{j})$, representing the experience’s core features, while the value contains the associated experiential data.
@@
$q_{u}=\text{BGE-M3}(u),~q_{\mathcal{V}}=\frac{1}{N}\sum^{N}_{i=1}\text{ImageEncoder}(\bar{v}_{i})$ (2)
where $q_{u}$ is generated by the BGE-M3 text encoder (46-chen2024bge), and $q_{\mathcal{V}}$ is produced by the image encoder of the MLLM. The stored key vectors $k_{u}^{j}$ and $k^{j}_{\mathcal{V}}$ for each Experience $E_{j}$ are computed analogously as $q_{u}$ and $q_{\mathcal{V}}$.
With the query vectors established, we measure their similarity to the key vectors of each stored experience $E_{j}$ by computing unimodal relevance scores for text $S_{u}$ and vision $S_{v}$ using cosine similarity:
@@
where $\alpha\in[0,1]$ balances the contribution of each modality. Ultimately, all experiences are ranked in descending order based on their MRS. We then filter this ranked list, selecting the top-$k$ experiences whose scores exceed a relevance threshold $\tau$. This procedure yields the final sets of retrieved experiences including the modality experience set $\{\mathcal{S}_{i}^{\text{ME}}\}^{k}_{i=1}$ and the semantic experience set $\{\mathcal{S}_{i}^{\text{SE}}\}^{k}_{i=1}$.
Experience-driven Stance Prediction.
@@
Following the retrieval of relevant past experiences, our ReMoD agent $A_{R}$ proceeds to generate initial, intuitive stance hypotheses. We formulate a comprehensive prompt that integrates two key components: (1) text information $u$ and visual information $\mathcal{V}$; (2) the retrieved experience sets $\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1}$ and $\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}$, which serve as in-context exemplars. This prompt instructs $A_{R}$ to perform a preliminary analysis from three distinct perspectives: text-only, vision-only, and combined multimodal. For each perspective, the agent generates an initial stance prediction $\hat{y}$ and a corresponding rationale $r$. This process is formally expressed as:
$(\hat{y}_{c},r_{c})=A_{R}(\text{prompt}(c,\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1},\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}))$ (5)
@@
where $c\in\{u,\mathcal{V},(u,\mathcal{V})\}$ denotes the context of evaluation: the text informationuu, the visual information $\mathcal{V}$ or the full multimodal pair $(u,\mathcal{V})$. These generated pairs $(\hat{y}_{c},r_{c})$ serve as essential intermediate reasoning traces that are scrutinized and refined during the subsequent deliberate reflective reasoning stage.
Deliberate Reflective Reasoning

### v2

# MIND Your Reasoning: A Meta-Cognitive Intuitive-Reflective Network for Dual-Reasoning in Multimodal Stance Detection

@@ Abstract
Multimodal Stance Detection (MSD) is a crucial task for understanding public opinion on social media. Existing methods predominantly operate by learning to fuse modalities. They lack an explicit reasoning process to discern how inter-modal dynamics, such as irony or conflict, collectively shape the user’s final stance, leading to frequent misjudgments. To address this, we advocate for a paradigm shift from learning to fuse to learning to reason. We introduce MIND, a Meta-cognitive Intuitive-reflective Network for Dual-reasoning. Inspired by the dual-process theory of human cognition, MIND operationalizes a self-improving loop. It first generates a rapid, intuitive hypothesis by querying evolving Modality and Semantic Experience Pools. Subsequently, a meta-cognitive reflective stage uses Modality-CoT and Semantic-CoT to scrutinize this initial judgment, distill superior adaptive strategies, and evolve the experience pools themselves. These dual experience structures are continuously refined during training and recalled at inference to guide robust and context-aware stance decisions. Extensive experiments on the MMSD benchmark demonstrate that our MIND significantly outperforms most baseline models and exhibits strong generalization.
MIND Your Reasoning: A Meta-Cognitive Intuitive-Reflective Network for Dual-Reasoning in Multimodal Stance Detection
Bingbing Wang1,2, Zhengda Jin1, Bin Liang3, Wenjie Li2, Jing Li2*, Ruifeng Xu1*, Min Zhang1, 1Harbin Institute of Technology, Shenzhen, China, 2The Hong Kong Polytechnic University, Hong Kong, China, 3The Chinese University of Hong Kong, Hong Kong, China
1 Introduction
@@
The proliferation of social platforms supporting multimodal content, such as text or images Dar et al. (2025), has made Multimodal Stance Detection (MSD) a critical task for understanding public opinion Liang et al. (2024); Niu et al. (2024). MSD aims to identify the expressed attitude toward a specific target by integrating both textual and visual information, enabling a more comprehensive analysis Liang et al. (2022). This task applies widely to political discourse analysis, misinformation detection, and public health communication, making it a crucial tool for decoding complex online interactions Niu et al. (2024); Pangtey et al. (2025).
Figure 1: Two examples contrasting existing methods (learning to fuse) in purple with our method (learning to reason) in pink.
While recent methods have advanced from static fusion to dynamic mechanisms like cross-attention, they remain fundamentally constrained by a paradigm of learning to fuse Liang et al. (2024); Zhang et al. (2025). This paradigm excels at weighing feature importance but lacks the explicit, high-level reasoning process needed to interpret inter-modal stance expressions, especially in semantic traps where seemingly consistent modalities convey a contradictory stance through nuanced cues like irony or cultural context. These traps manifest in at least two critical ways. First, models are easily deceived by cross-modal irony, as seen in Figure 1. A fusion-based model would likely interpret the positive text and the image objects as evidence for a FAVOR stance. It fails because it cannot perform the crucial reasoning step: recognizing that the text’s "best friends" sarcastically refers to the image’s "coffins," a grim symbol requiring world knowledge to interpret as a biting critique. Second, the challenge extends to intra-modal logical fallacies. A text can deceive a model focused on keywords like "sues" and "block" to conclude an AGAINST stance, completely missing that the "What if…" framing makes it a neutral COMMENT. In both scenarios, the failure stems not from an inability to weigh features, but from the absence of a high-level reasoning process to interpret non-literal intent and logical structure.
To bridge this gap, we advocate for a paradigm shift from learning to fuse to learning to reason. We introduce and pioneer MIND, a Meta-cognitive Intuitive-reflective Network for Dual-reasoning that realizes a more human-like, adaptive reasoning process for a complex multimodal stance detection. Inspired by the dual-process theory, MIND synergizes experience-driven intuitive reasoning with meta-cognitive reflective reasoning. First, an intuitive stage queries the Modality (MEP) and Semantic (SEP) Experience Pools to form an initial hypothesis from past reasoning patterns. This hypothesis is then scrutinized by a meta-cognitive reflective stage, where a Modality Chain of Thought (Modality-CoT) resolves inter-modal dynamics and formulates adaptive fusion strategies, while a complementary Semantic-CoT ensures a robust final decision. Extensive experiments on the public MMSD benchmark demonstrate that MIND significantly outperforms baseline models and exhibits strong generalization, validating the effectiveness of our meta-cognitive approach. Our main contributions are as follows:
  • We pioneer a paradigm shift for MSD from learning to fuse to learning to reason. Our dual-reasoning framework is the first to explicitly model a meta-cognitive process for resolving complex stance expressions.
  • We operationalize this paradigm with a novel self-improving meta-cognitive loop, where dual Experience Pools store and evolve explicit reasoning strategies that are then refined by a reflective Chain-of-Thought process.
@@
  • Experiments conducted on the public dataset demonstrate the superiority of our MIND, which significantly outperforms most baselines and exhibits strong generalization.
Figure 2: Overflow of our MIND framework for multimodal stance detection.
2 Related Works
2.1 LLM-based Stance Detection
@@
The design of modern stance detection frameworks has been significantly influenced by LLMs. Early approaches focused on harnessing their zero-shot and few-shot capabilities, using direct prompting and Chain-of-Thought (CoT) to elicit stance predictions without labeled data or fine-tuning Zhang et al. (2023, 2022). Other works have pursued fine-tuning by fusing retrieved background knowledge with input pairs to improve learning through backpropagation Li et al. (2023). Subsequent research has sought to further unleash LLM reasoning. Innovations include integrating multi-perspective insights via CoT and prompt-tuning Ding et al. (2024), formalizing reasoning with first-order logic rules for enhanced interpretability Zhang et al. (2024), and architecting multi-agent systems where LLMs collaborate for detection Lan et al. (2024); Wang et al. (2024). However, these methods are primarily developed for pure textual stance detection and have not yet explored the integration of multimodal information such as images.
2.2 Multimodal Stance Detection
@@
The proliferation of social media has driven the rise of multimodal content (text, image) in forms like memes and reels Dar et al. (2025); Bansal et al. (2024), which convey implicit opinions through diverse signals, posing challenges for stance detection. Modeling semantically heterogeneous cross-modal interactions remains difficult for traditional models, whereas transformer architectures and LLMs excel here via self-attention and transfer learning. Recent approaches improved stance classification through multimodal fusion. For example, Niu et al. (2024) proposed MmMtCSD and MLLM-SD for text-image pairs in multi-turn dialogues. Kuo et al. (2024) fused textual, visual, and social signals in political discourse using S-BERT, BEiT, and GraphSAGE. Xie et al. (2021) enhanced fake news detection via sentence-guided visual attention. Khiabani and Zubiaga (2023) and Barel et al. (2025) addressed few-shot or noisy settings with graph-based cues. Yao et al. (2023) extended MSD to fact-checking via web evidence integration. However, these approaches typically employ rigid fusion strategies, overlooking the dynamic interplay and context-dependent reliability of modalities, which makes them vulnerable to noisy or contradictory information.
3 Method
@@
This section introduces our novel MIND framework for MSD in detail. Drawing inspiration from the dual-process theory of human cognition Kahneman (2011), a psychological model that posits two distinct modes of thought. One is an intuitive, fast ‘System 1’ that relies on heuristics and prior experiences to make rapid judgments. The other is a deliberate, slow ‘System 2’ that engages in analytical reasoning to refine and validate these initial assessments. Our MIND framework operationalizes this cognitive paradigm for MSD.
@@
Given a text-image pair $(u,v)$, the aim of MIND is to identify the stance label $y$, towards a specific target $t$. As depicted in Figure 2, our MIND consists of three primary components: 1) Multimodal Knowledge Perception, which extracts comprehensive multimodal information and captures relevant external knowledge. 2) Experience-Driven Intuitive Reasoning (System 1), which queries the experience pools to form an initial stance hypothesis; and 3) Meta-cognitive Reflective Reasoning (System 2), which employs complementary reasoning chains to scrutinize the initial hypothesis and continuously refine the dual experience pools.
3.1 Multimodal Knowledge Perception
@@
3.1.1 Textual Knowledge Attainment
For the text input $u$, we perform a three-step knowledge extraction process. First, we employ a Multimodal Large Language Model (MLLM) with strategically designed prompts to identify a set of key semantic entities, denoted as $\mathcal{E}_{u}$. To ground these entities in external world knowledge, we retrieve corresponding background information for each entity $e\in\mathcal{E}_{u}$ from the Wikipedia. As the raw retrieved knowledge is often verbose and contains noisy information, we leverage the MLLM again to distill it into a concise, semantically-rich summary. This final summary, denoted as the text knowledge $\mathcal{K}_{u}$, serves as the knowledge for the text modality.
3.1.2 Visual Information Grounding
For the image input $v$, we capture both high-level semantic concepts and low-level visual information. First, to extract semantic concepts, an MLLM generates an image caption, identifies a set of visual entities $\mathcal{E}_{v}$ from the caption. These entities are then used to query Wikipedia, and the resulting information is summarized into a condensed visual knowledge $\mathcal{K}_{v}$. However, knowledge alone is insufficient to capture the visual manifestation of these concepts. To address this, we ground the semantic entities in the image’s pixel space. We leverage the identified entity set $\mathcal{E}_{v}$ to guide the open-vocabulary segmentation model SAM Ravi et al. (2024) in precisely segmenting the image regions corresponding to each entity, yielding a set of entity-aligned visual information $\mathcal{V}$ where each element is a segmented image directly showcasing the visual appearance of a specific semantic entity.
@@ 3.1.2 Visual Information Grounding
3.2 Experience-Driven Intuitive Reasoning
This stage simulates rapid, intuitive judgment by leveraging experience pools accumulated during training to quickly formulate an initial stance hypothesis. The process consists of two main steps: modal synergistic experience retrieval and experience-driven stance prediction.
3.2.1 Modal Synergistic Experience Retrieval
To retrieve relevant past experiences, we introduce a modal synergistic retrieval mechanism. Unlike conventional text-only systems, our approach leverages both text and visual information to query the modality and semantic experience pool. Each pool stores a collection of experiences $\{E_{j}\}$, where each experience is a key-value pair. The key is a pre-computed bimodal embedding $(k_{u}^{j},k_{\mathcal{V}}^{j})$, representing the experience’s core features, while the value contains the associated experiential data.
@@ 3.2.1 Modal Synergistic Experience Retrieval
$q_{u}=\text{Enc}(u),~q_{\mathcal{V}}=\frac{1}{N}\sum^{N}_{i=1}\text{Enc}(\bar{v}_{i})$ (2)
where $q_{u}$ and $q_{\mathcal{V}}$ are produced by the BGE-VL-CLIP Enc Zhou et al. (2025). The stored key vectors $k_{u}^{j}$ and $k^{j}_{\mathcal{V}}$ for each Experience $E_{j}$ are computed analogously as $q_{u}$ and $q_{\mathcal{V}}$. With the query vectors established, we measure their similarity to the key vectors of each stored experience $E_{j}$ by computing unimodal relevance scores for text $S_{u}$ and vision $S_{v}$ using cosine similarity:
@@
where $\alpha\in[0,1]$ balances the weight of each modality for retrieval. Ultimately, experiences are ranked in descending order based on their MRS. The top-$k$ experiences with scores excedding a relevance threshold $\tau$ are then selected, yielding the modality experience set $\{\mathcal{S}_{i}^{\text{ME}}\}^{k}_{i=1}$ and the semantic experience set $\{\mathcal{S}_{i}^{\text{SE}}\}^{k}_{i=1}$.
3.2.2 Experience-driven Stance Prediction
@@
Following the retrieval of relevant past experiences, our MIND agent $A_{R}$ proceeds to generate initial, intuitive stance hypotheses. We formulate a comprehensive prompt $\mathcal{P}$ that integrates two key components: (1) text information $u$ and visual information $\mathcal{V}$; (2) the retrieved experience sets $\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1}$ and $\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}$, which serve as in-context exemplars. This prompt instructs $A_{R}$ to perform a preliminary analysis from three distinct perspectives: text-only, vision-only, and combined multimodal. For each perspective, the agent generates an initial stance prediction $\hat{y}$ and a corresponding rationale $r$. This process is formally expressed as:
$(\hat{y}_{c},r_{c})=A_{R}(\mathcal{P}(c,\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1},\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}))$ (5)
@@
where $c\in\{u,\mathcal{V},(u,\mathcal{V})\}$ denotes the context of evaluation: the text information $u$, the visual information $\mathcal{V}$ or the full multimodal pair $(u,\mathcal{V})$. These generated pairs $(\hat{y}_{c},r_{c})$ serve as essential intermediate reasoning traces that are scrutinized and refined during the subsequent reflective reasoning stage.
3.3 Meta-cognitive Reflective Reasoning